OpenStudy (anonymous):

What is the solution to the following system? {3x+=3y=10 and -9x-9y=-30}

OpenStudy (compassionate):

Should we solve my substitution or elimination? What do you think?

OpenStudy (compassionate):

Also: 3x+=3y=10 lolwut

OpenStudy (anonymous):

3x+-3y=10*

OpenStudy (compassionate):

So, should we solve my elimination or substitution?

OpenStudy (anonymous):

substitution.

OpenStudy (compassionate):

3x+ - 3y=10 -9x-9y=-30 Take the first equation and solve it for y.

terenzreignz (terenzreignz):

Just checking... the +- sign is really how it's written? Sure that's not a typo?

OpenStudy (anonymous):

does y=10?

OpenStudy (compassionate):

Unfortunately I have to go.

OpenStudy (anonymous):

great..

terenzreignz (terenzreignz):

3x + - 3y Isn't it just 3x - 3y ?

OpenStudy (anonymous):

I'm writing it down exactly as it's showed on my page.

terenzreignz (terenzreignz):

Well, I'd say 3x + - 3y is 3x plus negative 3y which is just 3x minus 3y or 3x - 3y Unless I'm missing something.

OpenStudy (anonymous):

That's not what I'm worried about. I CANNOT solve this entire problem, I'm not asking you to simplify anything for me.

terenzreignz (terenzreignz):

It worries me. I'm simplifying it for me... :D So I need to know that you agree with me in that respect ^ So that we can continue on :)

OpenStudy (anonymous):

I agree.

terenzreignz (terenzreignz):

Okay. Substitution is just about the most straightforward way of solving systems of linear equations, but not always the most efficient. In fact, I'd say it gets highly inefficient the more variables you have. But now, you just have 2, so let's rewrite the equations for convenience... 3x - 3y = 10 -9x - 9y = 30 Shall we proceed?

OpenStudy (anonymous):

yes..

terenzreignz (terenzreignz):

Proceeding... The first step in solving systems of linear equations using substitution is pick a variable. To kickstart, let's pick x. What I want you to do is to isolate the x in the first equation. What I mean is I want you to rearrange that equation so that x stands alone on one side, and everything else in on the other.

OpenStudy (anonymous):

so x=10?

terenzreignz (terenzreignz):

Let's take a moment from solving... might I ask how you arrived at that solution?

OpenStudy (anonymous):

Darn. I have a correction for the original problem. You were right it was 3x+3y. But I subtracted 3y from each side. Then divided by 3.

terenzreignz (terenzreignz):

Okay... an error was discovered. Compensating.... Compensation complete. Rewriting system according to most recent update. 3x + 3y = 10 -9x - 9y = -30 Is this correct?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Foreseeing a possible complication. Ignoring. As usual, the first step is to isolate x in the first equation. Please do this now.

OpenStudy (anonymous):

yes yes I did. x=10

terenzreignz (terenzreignz):

This is impossible. If you are confused about isolating variables, please read through this sample. It takes you through the process. Remember that the main goal is simply to have x on one side of the equation and everything else on the other. \[\large 3x-12y=9\]Adding 12y to both sides yields \[\large 3x = 9 + 12y\]Dividing both sides by 3, yields... \[\large \frac{3x}{3}=\frac{9+12y}{3}\]\[\huge x=3+4y\]And thus x is isolated. Requesting that you process the first equation in a similar fashion.

mathslover (mathslover):

\[\large{\color{blue}{\textbf{WELCOME TO OPENSTUDY}}}\] @kickinmelvin